a particle is moving in x-y plane such that vx=4+4t &vy=4t/if the position of the particle is (1,2).then the equation of trajectory will be ?

Grade:12

2 Answers

Anil Pannikar AskiitiansExpert-IITB
13 years ago

 

Dear Lav Kumar,

equate v from 2 eq. and put value of x and y,  u will get t = -2

substitute t in 4+4t/x  = 4t y

so y = 2x is trajectory.

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Anil Pannikar

IIT Bombay

vikas askiitian expert
13 years ago

Vx=dx/dt

Vy=dy/dt

dy/dx=1/1+t

y=(t/1+t )x is the required equation

at (1,2) t=-2

 so the equation is y=2x